8) A store manager wishes to display 10 different brands of dog food in a row. In how many different ways can this be done? How many such arrangements are possible if he can use only 7 of the 10 different dog food types?
5) A store has 3 colors of appliances, white, grey, and black. The types of appliances are listed below. If an appliance is randomly selected, compute the following probabilities listed below. Blackta 5 White Grey 29 Refrigerator 10 Washer Dryer 14 12 16 13 40 a. The appliance is a refrigerator and grey. b. The appliance is white given that it is a washer. c. The appliance is a dryer or black.
Answer #1

8)

10 differnent brands of dog food can be arranged in 10! ways = 3628800 ways

if 7 of 10 different brands are to be used ,then this can done in = permutation(10,7) = 10!/3! = 604800 ways

5)

a)P(refrigeartor and grey) = 14/100 = 0.14

b) P(white | washer)=P(white and washer)/P(washer) = 11/31

c) P(dryer or black) = (13+16+11+8+5)/100 = 0.53